A solid spherical ball of mass 2.6 kg and radius 0.09 m rolls along a smooth, level surface with a speed of 2 m/s.  The ball rolls up an inclined plane.  How far up, vertical height, the plane does the ball go?  Express your answer as the vertical height from the level surface.
Answer #1

conserving energy,

0.5 mv^2 + 0.5Iw^2=mgh

or 0.5*mv^2 + 0.5*0.4*mv^2=mgh (I=0.4 mR^2 and v=wr)

or 0.5*1.4*mv^2=mgh

or 0.7*2^2=9.81*h

or h=0.285 m